OpenStudy (help_people):

A student is assessing the correlation between the average number of hours of internet browsing and average score on a math test for the students of a class. The table below shows the data: Number of hours of internet browsing (x) 0 0.5 1 1.5 2 2.5 3 3.5 4 4.5 Score on math test (y) 95 94 93 92 91 90 89 87 86 85

OpenStudy (help_people):

Part A: Is there any correlation between the average number of hours of internet browsing and average score on the math test for the students of the class? Justify your answer. Part B: Write a function which best fits the data. Part C: What does the slope and y-intercept of the plot indicate?

OpenStudy (welshfella):

what do you think? as the number of hours on the internet increase what happens to the math score?

OpenStudy (help_people):

decrease?

OpenStudy (welshfella):

yes - and also decrease very much in line

OpenStudy (welshfella):

for every 0.5 hour on the internet the math score goes down 1 point There is a definite negative correlation here.

OpenStudy (welshfella):

there is just one exception to this rule - between 3 and 3.5 hours

OpenStudy (welshfella):

where the mark drops by 2

OpenStudy (help_people):

ok?

OpenStudy (help_people):

so what would i say for pt a?

OpenStudy (welshfella):

Just summarise the above - there's a definite negative correlation born out by the data.

OpenStudy (welshfella):

for part B the function will be of the form y = mx + b y = mark and x = hours on internet where m is the slope of the graph of the function and b = value of the marks when the hours = 0, m will be negative

OpenStudy (welshfella):

for every rise of 0.5 in hours mark decreases by 1 so m = -1/0.5 = ? and b = 95 ( the mark when x = 0)

OpenStudy (welshfella):

can you figure out m now? -1 / 0.5 = ?

OpenStudy (help_people):

hold on srry

OpenStudy (help_people):

-2

OpenStudy (help_people):

@welshfella

OpenStudy (welshfella):

right so our function is y - -2x + 95 and the graph will look like

OpenStudy (help_people):

whats the -- stand for

OpenStudy (help_people):

or is the function y= -2x + 95?

OpenStudy (welshfella):